Find the nth term of the following sequences:

a.) 5, 10, 15, 20, …

b.) 22, 18, 14, 10, …


1
Expert's answer
2021-03-12T12:23:18-0500

a) This is an arithmetic sequence with "a_1=5,\;d=5."

"a_n=a_1+(n-1)d=5+5(n-1)=5n."


b) This is an arithmetic sequence with "a_1=22,\;d=-4."

"a_n=a_1+(n-1)d=22-4(n-1)=26-4n."